Okay, been messing with the problem for quite some time. Please help with set up, I seem to be missing something key to the problem:
How old is Blanche today if the square of her age 3 years from now, minus 10 times her age 2 years ago is 50. I keep thinking it must be something like :
(x+3)^2 - 10x-2=50
But it is not working out.

Answer by jim_thompson5910(35256)

the square of her age 3 years from now ---> (x+3)^2

10 times her age 2 years ago ----> 10(x-2)


-------------------------------------------------------

"How old is Blanche today if the square of her age 3 years from now, minus 10 times her age 2 years ago is 50"

turns into

(x+3)^2 - 10(x - 2) = 50

Solve for x

(x+3)^2 - 10(x - 2) = 50

x^2 + 6x + 9 - 10(x - 2) = 50

x^2 + 6x + 9 - 10x + 20 = 50

x^2 + 6x + 9 - 10x + 20 - 50 = 0

x^2-4x-21 = 0

(x-7)(x+3) = 0

x-7 = 0 or x+3 = 0

x = 7 or x = -3

Ignore the negative answer.

So the only solution is x = 7

This means that Blanche is 7 years old currently.